OpenStudy (anonymous):

If C is the triangle with vertices z=0, z=-4, and z=3i, what is the length of C? Please help!

OpenStudy (anonymous):

bfore clearing u r problem do u know what type of math is this??

OpenStudy (anonymous):

 is the picture, I'm pretty sure This is a complex calculus problem, and a small part of a larger question I am working on

OpenStudy (anonymous):

(it didn't come out very well, but thats a triangle from -4 to 3i to zero)

OpenStudy (anonymous):

u'll find difficulties as some what it is unrognized

OpenStudy (anonymous):

What do you mean by length of triangle??

OpenStudy (anonymous):

Sorry- Length of C, ie, the contour of the triangle

OpenStudy (anonymous):

@feldy90 go to this link! - http://www.webmath.com/

OpenStudy (anonymous):

sorry, I have never heard of contour of triangle. Can you give a better term for it??

OpenStudy (anonymous):

The entire question is somewhat more complex than this... That is, show \[| \int\limits_{c}^{} e^z-zdz|\le60\] where C is the triangle with vertices z=0, z=-4, z=3i And the method I am using is \[|\int\limits_{c}^{}e^z-zdz|\le \max(z \in C)|e^z-z|.\int\limits_{c}^{}|dz|\] And it is this final dz term which is the length of the contour. This is what my current question is about... Hope this clears anything up, and someone is still able to help?

OpenStudy (anonymous):

@dumbcow you have helped me quite a lot with various types of calculus before... Any insight to offer on this one? :)

OpenStudy (anonymous):

@Rohangrr thanks for the link, but I think my problem is slightly advanced for that site...

OpenStudy (anonymous):

no short it up! @feldy90

OpenStudy (dumbcow):

hmm don't know a lot regarding complex analysis...let me see if i study up a bit or find something to help you

OpenStudy (anonymous):

@feldy90 , is this your question \[\int\limits_{c}^{} \left| e ^{z} -z \right|dz \le60\] Because the position of modulus in your question is not clear.

OpenStudy (anonymous):

No, the modulus is indicated as being outside the integral, ie\[\left| \int\limits_{c}^{} e^z-zdz \right| \le 60\] Also, the z term is actually the conjugate of z (z-bar, if you like)

OpenStudy (anonymous):

And \[\left| \int\limits_{c}^{} e^z-zdz \right|\le \left(\begin{matrix}\max \\ z \in C\end{matrix}\right)\left| e^z-z \right|*\int\limits_{c}^{}\left| dz \right|\] Since this is true, I must compute the right hand side of this expression, and show that it is less than 60...

OpenStudy (anonymous):

@feldy90 , see this link --> http://mathhelpforum.com/calculus/77258-inequality-absolute-value-complex-integral.html This can help you I suppose.

OpenStudy (anonymous):

Answer = -2 + 2i interpreted Formula: (1+i)^3

OpenStudy (anonymous):

@feldy90

OpenStudy (anonymous):

@feldy90 , I got the thing you are looking for. See here --> http://www.math.ust.hk/~maykwok/courses/ma304/06_07/Complex_4.pdf

OpenStudy (anonymous):

Remember, with complex variables, keep like terms involving i together.... 1+i evaluates to 1+i Raising 1+i to the exponent 3 is the same as multiplying 1+i by itself 3 times. Multiply 1+i by 1+i Multiplying 1+i by 1+i is a classic Algebra problem. Here, you are trying to multiply two binomials together (two expressions that each contain two terms). Your book might call this finding the "Product of Two Binomials". To work this problem, we'll use the "F.O.I.L." method. F.O.I.L. stands for First, Outer, Inner, Last. First, we'll multiply the two First terms, the 1 and 1 together. 1 1 × 1 = 1 Second, we'll multiply the two Outer terms, the 1 and i together. Multiply 1 and i Multiply 1 and i The i just gets copied along. i 1 × i = i Third, we'll multiply the two Inner terms, the i and 1 together. Multiply i and 1 Multiply i and 1 The i just gets copied along. The answer is i i i × 1 = i i combines with i to give 2i Lastly, we'll multiply the two Last terms, the i and i together. Multiply i and i Multiply the i and i Multiply i and i Combine the i and i by adding the exponents, and keeping the i, to get The answer is i × i = Multiply by 1+i Multiply by 1+i we multiply by each term in 1+i term by term. This is the distributive property of multiplication. 1 1 × 1 = 1 Multiply 1 and i Multiply 1 and i The i just gets copied along. i 1 × i = i Multiply 2i and 1 Multiply i and 1 The i just gets copied along. The answer is i i 2i × 1 = 2i i combines with 2i to give 3i Multiply 2i and i Multiply the i and i Multiply i and i Combine the i and i by adding the exponents, and keeping the i, to get The answer is 2i × i = Multiply and 1 Multiply and 1 The just gets copied along. The answer is × 1 = combines with to give Multiply and i Multiply the and i Multiply and i Combine the and i by adding the exponents, and keeping the i, to get The answer is × i = (1+i)^3 evaluates to The final answer (almost!) is Now, let's simplify the i's to get our final answer: The i in 3i cannot be simplified, so just leave it as is. The i2 in can be simplified to -1. So, the becomes -3 . 1 combines with -3 to give -2 The i3 in can be simplified to -i. So, the becomes -i . 3i combines with -i to give 2i The final answer is -2+2i

OpenStudy (anonymous):

@Rohangrr This is taken from the webmath site you recommended, yes? It's not the addition or subtraction or multiplication of complex numbers which is my problem. I am wondering which vertices to add in the case of the triangle above to find my dz term To me, all I should be doing is summing the vertices, ie: 4 + 3i + sqrt(4^2 + 3i^2) (from pythagoras). This does not give me a reasonable answer, thus my asking for help from openstudy. Thank you, but I think we misunderstand each other...